OpenStudy (anonymous):

Please help me with this math. Determine the equation of the asymptotes g(x)=2x^2+1/x+5

OpenStudy (er.mohd.amir):

when fuction become infinite then x=ans is asymptotes

OpenStudy (anonymous):

Can you help me solve it?

OpenStudy (er.mohd.amir):

yes

OpenStudy (anonymous):

Ok so how do we start?

OpenStudy (er.mohd.amir):

if a/0=? where a is any number

OpenStudy (anonymous):

Ok

OpenStudy (er.mohd.amir):

so if x+5=0 then g(x) become infinity

OpenStudy (anonymous):

Ohh

OpenStudy (er.mohd.amir):

so x+5=0 is ans

OpenStudy (anonymous):

Thank you so much!